### v1

@@ 4.5   Global importance of IV surface features
where $\mathcal{C}_{j}$, the contribution of feature $j$ to the total risk reduction, is
$\mathcal{C}_{j}=\sum_{\mathcal{X}\subseteq\{\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R}\}\setminus\{j\}}\frac{|\cal{X}|!(5-|\cal{X}|)!}{6!}\left[\rho\left(\xi_{T}^{\tilde{\delta}_{\theta}(\cal{X})}\right)-\rho\left(\xi_{T}^{\tilde{\delta}_{\theta}(\cal{X},j)}\right)\right].$
Exact contributions $\mathcal{C}_{j}$ cannot be negative, although their estimates sometimes are. Negative values may occur for the contributions we present that are evaluated out-of-sample.
Our numerical experiments involve training RL agents using four penalty functions: CVaR95%, CVaR99%, MSE, and SMSE. We analyze the global importance of these state variables across different moneyness and maturities to comprehend their impact on the performance of the RL agents. The global importance of state variables is normalized by the risk reduction achieved by the respective RL agent to present contributions in the same order of magnitude: the relative global importance is
@@
Figure 6: Normalized global importance of features when hedging a European call options with a maturity $N=63$ days, across various moneyness levels.
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Each panel illustrates the Shapley values for all state variables ($(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ and different moneyness: OTM, ATM, and ITM. These results are shown for the four RL agents: RL-CVaR95%, RL-CVaR99%, RL-MSE, and RL-SMSE. The Shapley values are normalized by the risk reduction achieved by the respective agent according to Equation (9).
@@
Figure 7: Global importance of features when hedging a European ATM call options, across various maturities.
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Each panel illustrates the Shapley values for all features under different maturities: 21, 63, and 126 business days. These results are shown for the four RL agents: CVaR95%, CVaR99%, MSE, and SMSE. The Shapley values are normalized by the risk reduction achieved by the respective agent according to Equation (9).
4.6   Backtesting
In this section, we benchmark our approach using a historical path of the JIVR model spanning from January 5, 1996, to December 31, 2020, to assess the effectiveness of RL agents. Unlike our above experiments, which use simulated paths to test hedging strategies, this experiment examines their performance based on the historical series $\{R_{t},\beta_{t}\}$. Specifically, we evaluate the hedging performance considering a new European ATM call option with a maturity of 63 days every 21 business days along this historical path. The option prices, serving as initial hedging portfolio values, are determined with the prevailing IV surface on the day the hedge is initiated.
  • Results are computed based on the observed P&L from hedging 296 synthetic ATM Call options under real market conditions observed from May 1, 1996, to December 31, 2020. A new option is considered every 21 business days. Agents are trained under the reduced state space $(\delta_{t},\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ according to the conditions outlined in Section 4.3.1.
@@ 5   Conclusion
This study introduces a novel deep hedging framework that integrates forward-looking volatility information through a functional representation of the IV surface, combined with backward-looking conventional features. Our implementation employs deep policy gradient methods and utilizes a unique neural network architecture consisting of LSTM cells and FFNN layers to enhance training efficiency. Additionally, the architecture incorporates a budget constraint mitigating the incentive to gamble and enabling the agent to learn hedging strategies that are more effective for risk management.
@@
The global importance analysis of IV features confirms the significant enhancement of hedging performance in terms of risk reduction relative to the penalty functions. Our analysis underscores the critical importance of key factors such as the underlying asset return conditional variance process ($h_{R}$), the long-term ATM IV level ($\beta_{1}$) and the time-to-maturity slope ($\beta_{2}$). RL agents utilize both the historical variance process and market expectations of future volatility to adjust positions in the underlying asset.
Appendix A Details for the MSGD training approach
@@
[…]
@@
Appendix B Joint Implied Volatility and Return model
B.1   Daily implied volatility surface
@@
B.2   Joint Implied Volatility and Return
@@
the process $\{\epsilon_{t,R}\}^{T}_{t=0}$ is a sequence of iid standardized NIG random variables with parameters $\zeta_{R}$ and $\varphi_{R}$, [^10] The standard NIG random variable $\epsilon$ follows the probability density function with parameters $\zeta$ and $\varphi$: $f(x)=\frac{B_{1}\left(\sqrt{\frac{\varphi^{6}}{\varphi^{2}+\zeta^{2}}+(\varphi^{2}+\zeta^{2})\left(x+\frac{\varphi^{2}\zeta}{\varphi^{2}+\zeta^{2}}\right)^{2}}\right)}{\pi\sqrt{\frac{1}{\varphi^{2}+\zeta^{2}}+\frac{\varphi^{2}+\zeta^{2}}{\varphi^{6}}\left(x+\frac{\varphi^{2}\zeta}{\varphi^{2}+\zeta^{2}}\right)^{2}}}e^{\left(\frac{\varphi^{4}}{\varphi^{2}+\zeta^{2}}+\zeta\left(x+\frac{\varphi^{2}\zeta}{\varphi^{2}+\zeta^{2}}\right)\right),}$ where $B_{1}(\cdot)$ denotes the modified Bessel function of the second kind with index 1. The common $(\alpha,\beta,\delta,\mu)$-specification can be obtained by replacing $\beta$ and $\gamma$ ($\gamma=\sqrt{\alpha^{2}-\beta^{2}}$), with $\zeta$ and $\varphi$, respectively, and imposing a null mean and unit variance to express $\delta$ and $\mu$ interms of $\alpha$, $\beta$. and $\psi$ is their cumulant generating function. [^11] For $-\sqrt{\zeta^{2}+\varphi^{2}}-\zeta<z<\sqrt{\zeta^{2}+\varphi^{2}}-\zeta$, the cumulant generating function is given by $\psi(z)=\frac{\varphi^{2}}{\varphi^{2}+\zeta^{2}}\left(-\zeta z+\varphi^{2}-\varphi\sqrt{\varphi^{2}+\zeta^{2}-(\varphi+\zeta)^{2}}\right).$ Parameters of the excess return component of the model are thus $\Theta_{R}=(\lambda,\kappa_{R},\gamma_{R},a_{R},\omega_{R},\zeta_{R},\varphi_{R})$.
@@
Estimates of all JIVR model parameters are taken from Table 5 and Table 6 of François et al., (2023). In all experiments, the annualized continuously compounded risk-free rate and dividend yield are assumed to be constant with values set to $r=2.66\%$ and $q=1.77\%$, respectively.
Appendix C Benchmarks
C.1   Black-Scholes model
$\mbox{Call}_{t}=S_{t}\cdot\mbox{e}^{-q_{t}\tau_{t}}\Phi(d_{t})-K\cdot e^{-r_{t}\tau_{t}}\Phi(d_{t}-\sigma(M_{t},\tau_{t},\beta_{t})\sqrt{\tau_{t}})$
$\Delta_{t}^{BS}=\mbox{e}^{-q_{t}\tau_{t}}\Phi(d_{t}).$
C.2   Leland model
$\Delta_{t}^{L}=\mbox{e}^{-q_{t}\tau_{t}}\Phi\left(\tilde{d}_{t}\right),$
C.3   Smile-implied model
with
and
Appendix D Network fine-tuning
D.1   Bounded strategies - leverage constraint
@@
D.2   Network architecture selection
@@
The superiority of the RNN-FNN over the LSTM and the FFNN is evaluated based on the optimal value of each penalty function. Figure 10 illustrates the optimal values of each penalty function for each architecture, normalized by the estimated value obtained with the BS delta. Notably, the RNN-FNN setup considered in this paper significantly outperforms both the benchmark and other architectures. Conversely, the FFNN does not surpass the benchmark for the MSE, and the LSTM exhibits almost the same performance as the benchmark.
@@
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ according to the conditions outlined in Section 4.3.1. The setup for the different networks follows the architecture described in Section 2.2.1, with $L_{1}=0$ and $L_{2}=4$ for the FFNN, $L_{1}=4$ and $L_{2}=0$ for the LSTM, and $L_{1}=2$ and $L_{2}=2$ for RNN-FNN. Results show optimal values obtained from agents trained under CVaR95%, CVaR99%, MSE, and SMSE for the three networks. These values are normalized by the estimated values of each penalty function obtained with the BS delta.
A second test to demonstrate the superiority of our architecture involves computing the optimal values of the penalty functions over various clusters of paths. The objective is to isolate the impact of different state of the economy on performance and assess the robustness of our architecture. LABEL:fig:networks_state_economy displays the optimal values of each architecture normalized by the estimated values obtained by the BS delta for all penalty functions. Notably, RNN-FNN agents outperform both the benchmark and other architectures across all penalty functions, regardless of the economic conditions under which the simulations were conducted. Conversely, FFNN and LSTM agents fail to outperform the benchmark across all states of the economy when the agents are trained under the CVaR95%, MSE and SMSE penalty functions, as shown in the top-left, bottom-left and bottom-right panels, respectively.
@@
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ according to the conditions outlined in Section 4.3.1. Each panel illustrates the values obtained by different architectures for the following hedging metrics: Avg P&L, CVaR95%, MSE, and SMSE. These metrics are normalized by the estimated values of each penalty function obtained with the BS delta. The setup for the different networks follows the architecture described in Section 2.2.1, with $L_{1}=0$ and $L_{2}=4$ for the FFNN, $L_{1}=4$ and $L_{2}=0$ for the LSTM, and $L_{1}=2$ and $L_{2}=2$ for RNN-FNN.131313 The periods aim to approximate different states of the economy considering the time frames specified in Table 2.
D.3   Dropout parameter selection

### v2

@@ 4.5  Global importance of IV surface factors
where $\mathcal{C}_{j}$, the contribution of variable $j$ to the total risk reduction, is
[…]
Exact contributions $\mathcal{C}_{j}$ cannot be negative, although their estimates sometimes are. Negative values may occur for the contributions we present that are evaluated out-of-sample. Our numerical experiments involve training RL agents using four penalty functions: CVaR95%, CVaR99%, MSE, and SMSE. We analyze the global importance of these state variables across different moneyness and maturities to comprehend their impact on the performance of the RL agents. The global importance of state variables is normalized by the risk reduction achieved by the respective RL agent to present contributions in the same order of magnitude: the relative global importance is
@@
Figure 6: Normalized global importance of state variables when hedging a European call options with a maturity $N=63$ days, across various moneyness levels.
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Each panel illustrates the Shapley values for all state variables $(\{\beta_{t,i}\}^{5}_{i=1},h_{t,R})$ and different moneyness: OTM, ATM, and ITM. These results are shown for the four RL agents: RL-CVaR95%, RL-CVaR99%, RL-MSE, and RL-SMSE. The Shapley values are normalized by the risk reduction achieved by the respective agent according to Equation (9).
Figure 7: Global importance of state variables when hedging a European ATM call options, across various maturities.
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Each panel illustrates the Shapley values for all state variables under different maturities: 21, 63, and 126 business days. These results are shown for the four RL agents: CVaR95%, CVaR99%, MSE, and SMSE. The Shapley values are normalized by the risk reduction achieved by the respective agent according to Equation (9).
@@
5  Backtesting
We conduct out-of-sample backtesting experiments to assess the model’s performance on actual data. Specifically, we evaluate the effectiveness of our approach using traded option prices observed between December 31, 2020, and October 31, 2023, ensuring that the RL agents, trained on simulated data reflecting estimated dynamics from January 1996 to December 2020, are tested on realized market data they have never previously encountered.
All options enter the sample as soon as their remaining time-to-maturity is 63 business days, at which point a short position is taken and hedged.111111 To ensure liquidity and reliability throughout the hedging period, we retain only options that report non-zero values for best bid, trading volume, and open interest at the onset. These filters are based on the fields best_bid, volume, and open_interest from OptionMetrics. We include all options that meet the criteria within the test set period. All retained options are European calls that are around-the-money, defined as having an underlying price within $\pm 10\%$ of the strike. [^12] Because option prices are proportional to the initial value of the underlying asset in our framework, we normalize all prices such that the initial value of the underlying is \$100. This allows us to replicate conditions similar to those used to train the RL algorithms, without any loss of generality. This leaves us with 4,134 hedged contracts. The state variables are reconstructed from historical data using François et al., (2022) and François et al., (2023). In the presence of transaction costs, the state space is augmented to include the current portfolio position, resulting in $(\tau_{t},S_{t},\{\beta_{t,i}\}_{i=1}^{5},h_{t,R},\delta_{t})$, where $\delta_{t}$ denotes the number of shares held at time $t$.
We compare the performance of benchmark strategies—DH, SI, and DH-L (with transaction costs)—against our RL-based approaches. We also include a benchmark RL agent that excludes the IV surface from its input features, namely $\{\beta_{t,i}\}^{5}_{i=1}$. This restricted RL strategy exhibits weaker risk reduction in the simulation environment, as highlighted by the Shapley value decomposition from Section˜4.5 which reveals a positive contribution from the incorporation of IV coefficients. We evaluate whether this performance gap persists in the out-of-sample backtest. All strategies are assessed using three risk metrics: MSE, SMSE, and CVaR95%.
Figure 8: Backtest risk metrics on hedging errors, with and without transaction costs.
  • Backtests are conducted on 4,134 around-the-money call options, using actual market prices observed between December 31, 2020 and October 31, 2023.
As shown in Figure 8 the RL algorithm with IV information and the smile-implied delta hedging, both of which leverage the IV surface, deliver comparable performance in terms of MSE. They are closely followed by the RL algorithm with restricted information, while the practitioners’ delta hedging lags significantly behind. When transaction costs are introduced, both RL algorithms outperform all variants of delta hedging strategies. When focusing on tail risk measures—SMSE and CVaR95%—RL strategies consistently outperform traditional benchmarks. Moreover, incorporating the IV surface information yields additional performance gains.
Figure 9: Backtest hedging error distributions, without transaction costs.
  • The results are based on hedging 4134 around-the-money call options using real market prices observed between December 31, 2020 and October 31, 2023. No transaction costs are applied.
The superior performance of RL strategies over delta-based benchmarks is further illustrated in Figure 9. When the risk metric is the MSE, the goal is to center the hedging error around zero. As shown in Panel A, the RL strategy is more successful in achieving such objective than the practitioner’s delta hedge.
For asymmetric risk metrics such as SMSE and CVaR, the objective shifts toward limiting exposure to extreme losses. Panels B and C show that the RL strategies produce notably thinner right tails in the hedging error distributions, reflecting a more effective control of downside risk. A similar pattern is observed in panels D, E, and F, where RL strategies with the full market information also outperform their counterparts that exclude the IV surface information.
These findings suggest that RL-based hedging can serve as an effective risk management tool. However, the inclusion of relevant information from the implied volatility surface is required to unleash its full potential; its performance depends on the quality and completeness of the input data.
@@ 6  Conclusion
This study introduces a novel deep hedging framework that integrates forward-looking volatility information through a functional representation of the IV surface, combined with backward-looking conventional state variables. Our implementation employs deep policy gradient methods and utilizes a neural network architecture consisting of LSTM cells and FFNN layers to enhance training efficiency. Additionally, the architecture incorporates a budget constraint mitigating the incentive to gamble and enabling the agent to learn hedging strategies that are more effective for risk management.
@@
The global importance analysis of IV factors confirms the significant enhancement of hedging performance in terms of risk reduction relative to the penalty functions. Our analysis underscores the critical importance of key factors such as the underlying asset return conditional variance process ($h_{R}$), the long-term ATM IV level ($\beta_{1}$) and the time-to-maturity slope ($\beta_{2}$). RL agents utilize both the historical variance process and market expectations of future volatility to adjust positions in the underlying asset. Out-of-sample backtests underscore the critical role of IV surface information in unlocking the full potential of RL-based hedging strategies.
Appendix A Benchmarks
$\mbox{Call}_{t}=S_{t}\mbox{e}^{-q_{t}\tau_{t}}\Phi(d_{t})-Ke^{-r_{t}\tau_{t}}\Phi(d_{t}-\sigma(M_{t},\tau_{t},\beta_{t})\sqrt{\tau_{t}})$
$\Delta_{t}^{BS}=\mbox{e}^{-q_{t}\tau_{t}}\Phi(d_{t}).$
$\Delta_{t}^{L}=\mbox{e}^{-q_{t}\tau_{t}}\Phi\left(\tilde{d}_{t}\right),$
with
and
Supplementary material (not part of the paper)
Appendix B Details for the MSGD training approach
@@
where $\widehat{\mbox{VaR}}_{\alpha}(\mathbb{B}_{j})=\xi_{T,\lceil\alpha\cdot N_{\mbox{\scriptsize{batch}}}\rceil}^{\tilde{\delta}_{\theta_{j}}}$ is the estimation of the VaR obtained from the ordered sample $\{\xi_{T,[i]}^{\tilde{\delta}_{\theta_{j}}}\}_{i=1}^{N_{\mbox{\scriptsize{batch}}}}$ and $\lceil\cdot\rceil$ is the ceiling function. These empirical approximations are used to estimate the gradient of the penalty function required in Equation (4). [^13] In particular, the gradient of these estimations has analytical expressions for FFNN, LSTM networks and thus for RNN-FNN networks. Details about gradient of the empirical objective function are provided in Goodfellow et al., (2016).
@@
Appendix C Joint Implied Volatility and Return model
C.1  Daily implied volatility surface
@@
C.2  Joint Implied Volatility and Return
@@
the process $\{\epsilon_{t,R}\}^{T}_{t=0}$ is a sequence of iid standardized NIG random variables with parameters $\zeta_{R}$ and $\varphi_{R}$, [^14] The standard NIG random variable $\epsilon$ follows the probability density function with parameters $\zeta$ and $\varphi$: $f(x)=\frac{B_{1}\left(\sqrt{\frac{\varphi^{6}}{\varphi^{2}+\zeta^{2}}+(\varphi^{2}+\zeta^{2})\left(x+\frac{\varphi^{2}\zeta}{\varphi^{2}+\zeta^{2}}\right)^{2}}\right)}{\pi\sqrt{\frac{1}{\varphi^{2}+\zeta^{2}}+\frac{\varphi^{2}+\zeta^{2}}{\varphi^{6}}\left(x+\frac{\varphi^{2}\zeta}{\varphi^{2}+\zeta^{2}}\right)^{2}}}e^{\left(\frac{\varphi^{4}}{\varphi^{2}+\zeta^{2}}+\zeta\left(x+\frac{\varphi^{2}\zeta}{\varphi^{2}+\zeta^{2}}\right)\right),}$ where $B_{1}(\cdot)$ denotes the modified Bessel function of the second kind with index 1. The common $(\alpha,\beta,\delta,\mu)$-specification can be obtained by replacing $\beta$ and $\gamma$ ($\gamma=\sqrt{\alpha^{2}-\beta^{2}}$), with $\zeta$ and $\varphi$, respectively, and imposing a null mean and unit variance to express $\delta$ and $\mu$ interms of $\alpha$, $\beta$. and $\psi$ is their cumulant generating function. [^15] For $\sqrt{\zeta^{2}+\varphi^{2}}-\zeta<z<\sqrt{\zeta^{2}+\varphi^{2}}-\zeta$, the cumulant generating function is given by $\psi(z)=\frac{\varphi^{2}}{\varphi^{2}+\zeta^{2}}\left(-\zeta z+\varphi^{2}-\varphi\sqrt{\varphi^{2}+\zeta^{2}-(\varphi+\zeta)^{2}}\right).$ Parameters of the excess return component of the model are thus $\Theta_{R}=(\lambda,\kappa_{R},\gamma_{R},a_{R},\omega_{R},\zeta_{R},\varphi_{R})$.
@@
Estimates of all JIVR model parameters are taken from Table 5 and Table 6 of François et al., (2023).
Appendix D JIVR Model parameters
Table 4: Estimated Gaussian copula parameters
$\epsilon_{t,R}$ 1.000
$\epsilon_{t,1}$ -0.550 1.000
$\epsilon_{t,2}$ -0.690 0.140 1.000
$\epsilon_{t,3}$ 0.030 -0.030 -0.0100 1.000
$\epsilon_{t,4}$ -0.220 0.250 0.120 0.280 1.000
$\epsilon_{t,5}$ -0.340 0.170 0.370 0.130 -0.050 1.000
Table 5: JIVR model parameter estimates
Parameter $\beta_{1}$ $\beta_{2}$ $\beta_{3}$ $\beta_{4}$ $\beta_{5}$ S&P500
$\theta_{1}$ 0.996290 -0.013869 0.002841
$\theta_{2}$ 0.003669 0.877813 0.001300
$\theta_{3}$ -0.032640 0.997071 0.003722 -0.004198
$\theta_{4}$ 0.980269
$\theta_{5}$ -0.047789 0.986019
$\nu$ 0.089445
$\sigma\sqrt{252}$ 0.380279 0.052198 0.048641 0.051536
$\omega$ 0.267589 0.977291
$\kappa$ 0.838220 0.965751 0.974251 0.945377 0.980844 0.888977
$a$ 0.134152 0.098272 0.092646 0.102201 0.100502 0.056087
$\gamma$ -0.111813 -1.482862 0.096766 0.060558 -0.102996 2.507796
$\zeta$ 0.143760 0.852943 0.029109 -0.159051 0.092664 -0.641306
$\varphi$ 1.351070 1.538928 2.284780 1.449977 1.428477 2.039669
Appendix E Network fine-tuning
E.1  Bounded strategies - leverage constraint
@@
E.2  Network architecture selection
@@
The superiority of the RNN-FNN over the LSTM and the FFNN is evaluated based on the optimal value of each penalty function. Figure 11 illustrates the optimal values of each penalty function for each architecture, normalized by the estimated value obtained with DH. Notably, the RNN-FNN setup considered in this paper significantly outperforms both the benchmark and other architectures. Conversely, the FFNN does not surpass the benchmark for the MSE, and the LSTM exhibits almost the same performance as the benchmark.
@@
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ according to the conditions outlined in Section 4.3.1. The setup for the different networks follows the architecture described in Section 2.2.1, with $L_{1}=0$ and $L_{2}=4$ for the FFNN, $L_{1}=4$ and $L_{2}=0$ for the LSTM, and $L_{1}=2$ and $L_{2}=2$ for RNN-FNN. Results show optimal values obtained from agents trained under CVaR95%, CVaR99%, MSE, and SMSE for the three networks. These values are normalized by the estimated values of each penalty function obtained with DH.
A second test to demonstrate the superiority of our architecture involves computing the optimal values of the penalty functions over various clusters of paths. The objective is to isolate the impact of different state of the economy on performance and assess the robustness of our architecture. LABEL:fig:networks_state_economy displays the optimal values of each architecture normalized by the estimated values obtained by DH for all penalty functions. Notably, RNN-FNN agents outperform both the benchmark and other architectures across all penalty functions, regardless of the economic conditions under which the simulations were conducted. Conversely, FFNN and LSTM agents fail to outperform the benchmark across all states of the economy when the agents are trained under the CVaR95%, MSE and SMSE penalty functions, as shown in the top-left, bottom-left and bottom-right panels, respectively.
@@
  • Results are computed using 100,000 out-of-sample paths in the absence of transaction costs ($\kappa=0\%$). Agents are trained under the reduced state space $(\tau_{t},S_{t},\{\beta_{t,i}\}^{5}_{i=1},h_{t,R},\{h_{t,i}\}^{5}_{i=1})$ according to the conditions outlined in Section 4.3.1. Each panel illustrates the values obtained by different architectures for the following hedging metrics: Avg P&L, CVaR95%, MSE, and SMSE. These metrics are normalized by the estimated values of each penalty function obtained with DH. The setup for the different networks follows the architecture described in Section 2.2.1, with $L_{1}=0$ and $L_{2}=4$ for the FFNN, $L_{1}=4$ and $L_{2}=0$ for the LSTM, and $L_{1}=2$ and $L_{2}=2$ for RNN-FNN.171717 The periods aim to approximate different states of the economy considering the time frames specified in Table 2.
E.3  Dropout parameter selection